Solve the following equation, answer as a reduced, mixed number.6(x - 1) = 4(x - 2)
Answer by lenny460(1073)

6(x - 1) = 4(x - 2)
6 * x = 6x
6 * -1 = -6
6x - 6 = 4(x - 2)
4 * x = 4x
4 * -2 = -8
6x - 6 = 4x - 8
Subtract 4x from each side of the equation.
6x - 6 - 4x = 4x - 8 - 4x
Combine like terms:
6x - 4x = 2x
4x - 4x = 0
Therefore:
2x - 6 = -8
Add 6 to each side of the equation.
2x - 6 + 6 = -8 + 6
Combine like terms:
-6 + 6 = 0
-8 + 6 = -2
Therefore:
2x = -2
Divide each side by 2
2x/2 = -2/2
x = -1
The Answer:
x = -1


obuong3@aol.com